OpenStudy (anonymous):

area is 46sq.in and height is 20in a=1/2bh what is the base?

OpenStudy (chaise):

Substitute your values: a=1/2*b*h 46=1/2*b*20 46=b*10 b=46/10 b=4.6

OpenStudy (anonymous):

thank you